OpenStudy (anonymous):

One way to solve the inequality -│x - 2│ + 3 ≥ 0 can somebody show me how to do this

OpenStudy (anonymous):

-|x-2|>=-3 |x-2|>=3 now solve it???

OpenStudy (anonymous):

i dont no how to

myininaya (myininaya):

Don't forget multiplying by negative changes the direction of the inequality

OpenStudy (anonymous):

i told you that |x|=x when x>0 & |x|=-x when x<0 therefore x-2<=3 x<=5 &&& -(x-2)<=3 x-2>=3 x>=5 also what@myininaya said is right the last line of my previous comment is |x-3|<=3